OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 2x + 3y = 11 5x + 3y = 5 A. (-2, 5) B. (2, -5) C. (-2, -5) D. (2, 5)

OpenStudy (anonymous):

have you tried anything at all? ... or you don't know how to start it ? :)

OpenStudy (anonymous):

didn't know how too start and just dk how to do it

OpenStudy (anonymous):

ok, so you have: 2x + 3y = 11 5x + 3y = 5 let's start from the first equation: 2x + 3y = 11 we'll solve it for X 2x=11-3y now let's divide both sides by 2 and we'll have X alone in the left side: \[\Large \frac{2x}{2}=\frac{11}{2}-\frac{3y}{2}\] \[\Large x=\frac{11}{2}-\frac{3y}{2}\] now take the second equation... and substitute X \[\Large 5x+3y=5\] \[\Large 5\left(\frac{11}{2}-\frac{3y}{2}\right)+3y=5\] now multiply\simplify anything that you havet o.. and find y, then tell me what did you get. I'd be delighted to check it :) ...if you get stuck somewhere just let me know.

OpenStudy (anonymous):

(-2, 5)

OpenStudy (anonymous):

how did you get it.. ? :) would you care to show your work ?

OpenStudy (anonymous):

i showed my work on a sheet of paper is it right

OpenStudy (anonymous):

unfortunately I don't see it. ... O_O :)

OpenStudy (anonymous):

is my answer right

OpenStudy (anonymous):

yes it is... but I also would like to know if you know to solve it or not ?

OpenStudy (anonymous):

i do now

OpenStudy (anonymous):

ok.. I'm waiting.